OpenStudy (anonymous):

12 - Caitlin had $402 in her bank account. She withdrew $15 each week to pay for a swimming lesson. She now has $237. a. Write an equation that can be used to find the number of swimming lessons that she paid for. b. How many swimming lessons did she pay for? c. At the time she has $237, the cost of a lesson rose to $19. How many lessons can she pay for with her remaining $237?

OpenStudy (anonymous):

402 - 237 = 165 165/15 = 11 (402 - 237) ---------- 15 She paid for 15 swimming lessons 237 / 19 = 12.47 She can pay for 12 more lessons